OpenStudy (anonymous):

Line segment AB has endpoints (–3.8,–2.7) and (4.3,3.4). Line segment CD has endpoints (8.7,–2.8) and (14.8,5.3). Determine the length of each of these line segments. Then decide whether the two line segments are congruent, and justify your choice. Write your answers with one decimal place.Are the two line segments congruent?Length of line segment AB =Length of line segment CD =

OpenStudy (anonymous):

@Mehek14 @escamer

jimthompson5910 (jim_thompson5910):

Are you familiar with the distance formula?

jimthompson5910 (jim_thompson5910):

This formula \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\]

OpenStudy (anonymous):

not exactly

jimthompson5910 (jim_thompson5910):

ok so your first step is to subtract the corresponding x coordinates in this case, -3.8 and 4.3 I'm focusing on the points (–3.8,–2.7) and (4.3,3.4)

jimthompson5910 (jim_thompson5910):

what is -3.8 minus 4.3 equal to?

OpenStudy (anonymous):

-8.1

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now we square that result to get what?

OpenStudy (anonymous):

-65.1

jimthompson5910 (jim_thompson5910):

(-8.1)^2 = 65.61 keep in mind that squaring a negative produces a positive result

jimthompson5910 (jim_thompson5910):

since negative times negative = positive

jimthompson5910 (jim_thompson5910):

ok so we have that result. We need to keep it in mind for later now we move onto the y coordinates of (–3.8,–2.7) and (4.3,3.4) we do the same steps -2.7 minus 3.4 = ???

OpenStudy (anonymous):

-6.1

jimthompson5910 (jim_thompson5910):

square that to get what?

OpenStudy (anonymous):

-37.21

jimthompson5910 (jim_thompson5910):

again, remember when you square a negative, you get a positive

jimthompson5910 (jim_thompson5910):

(-6.1)^2 = (-6.1)*(-6.1) = +37.21

OpenStudy (anonymous):

so 37.21 and 65.1

jimthompson5910 (jim_thompson5910):

(-8.1)^2 = 65.61 idk how you're getting 65.1

jimthompson5910 (jim_thompson5910):

The two results after squaring are 65.61 (from the x terms) 37.21 (from the y terms)

OpenStudy (anonymous):

i was typing it wromg in the calculator. so AB= 65.61 CD=37.21 Are they congruent?

jimthompson5910 (jim_thompson5910):

we haven't moved to CD yet. We're still on AB

jimthompson5910 (jim_thompson5910):

we're not done with AB

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

ok so we have these two results: 65.61 and 37.21 next step is to add them up 65.61 + 37.21 = 102.82 then finally, the last step is to take the square root of that result \[\Large \sqrt{102.82} \approx 10.1400197\] so AB is approximately 10.1400197 units long

OpenStudy (anonymous):

Ok.

jimthompson5910 (jim_thompson5910):

Recap: Step 1) Subtract the x coordinates Step 2) Square the result from step 1 Step 3) Subtract the y coordinates Step 4) Square the result from step 3 Step 5) Add the results from steps 2 and 4 Step 6) Take the square root from the result of step 5

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

Tell me what you get for CD

OpenStudy (anonymous):

8.1 6.1 65.61 65.61 131.22 17218.68

OpenStudy (anonymous):

i mean 11.4

jimthompson5910 (jim_thompson5910):

let me check, one moment

jimthompson5910 (jim_thompson5910):

you have two copies of 65.61. Double check that part

OpenStudy (anonymous):

10.1

jimthompson5910 (jim_thompson5910):

Distance from C(8.7,–2.8) and D(14.8,5.3) Step 1) Subtract the x coordinates 8.7-14.8 = -6.1 Step 2) Square the result from step 1 (-6.1)^2 = 37.21 Step 3) Subtract the y coordinates 5.3 - (-2.8) = 5.3 + 2.8 = 8.1 Step 4) Square the result from step 3 8.1^2 = 65.61 Step 5) Add the results from steps 2 and 4 37.21 + 65.61 = 102.82 Step 6) Take the square root from the result of step 5 \(\Large \sqrt{102.82} \approx 10.1400197\) so we see that AB = CD because they are the same length. The distance from A to B, is the same as the distance from C to D

jimthompson5910 (jim_thompson5910):

btw, you can subtract the x coordinates in any order notice how 8.7 - 14.8 = -6.1 square that to get 37.21 and how 14.8 - 8.7 = 6.1 square that to get 37.21 The results after squaring the x difference are the same. So that's why you can subtract in any order. The same applies to y as well.

jimthompson5910 (jim_thompson5910):

hopefully all this makes sense?

OpenStudy (anonymous):

Yes thank you!!!

jimthompson5910 (jim_thompson5910):

ok great, you're welcome